# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ 3 Mirror descent
For every $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let $\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ be strictly convex, proper and differentiable and assume that the (sub)gradient $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\in L^{2}(\mu)$ exists. In this section, we are interested in analyzing the scheme (1) where the cost $\mathrm{d}$ is chosen as a Bregman divergence, i.e. $\mathrm{d}_{\phi_{\mu}}$ as defined in Definition 1. This corresponds to a mirror descent scheme in $\mathcal{P}_{2}(\mathbb{R}^{d})$:
@@
Iterates of MD.
In all that follows, we assume that the iterates (4) exist, which is true e.g. for a superlinear $\phi_{\mu_{k}}$, since the objective is a sum of linear functions and of the continuous $\phi_{\mu_{k}}$. In the previous section, we have seen that the second term in the proximal scheme (4) can be interpreted as a linearization of the functional $\mathcal{F}$ at $\mu_{k}$ for Wasserstein (sub-)differentiable functionals. Now define for all $\mathrm{T}\in L^{2}(\mu_{k})$, $\mathrm{J}(\mathrm{T})=\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})+\tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}$. Then, deriving the first order conditions of (4) as $\nabla\mathrm{J}(\mathrm{T}_{k+1})=0$, we obtain $\mu_{k}$-a.e.,
@@
Note that for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|^{2}_{L^{2}(\mu)}$, the update (5) translates as $\mathrm{T}_{k+1}=\mathrm{Id}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})$, and our scheme recovers Wasserstein gradient descent [32, 84]. This is analogous to mirror descent recovering gradient descent when the Bregman potential is chosen as the Euclidean squared norm in $\mathbb{R}^{d}$ [11]. We discuss in Section D.2 the continuous formulation of (4), showing it coincides with the gradient flow of the mirror Langevin [3, 119], the limit of the JKO scheme with Bregman groundcosts [97], Information Newton’s flows [115], or Sinkhorn’s flow [38] for specific choices of $\phi$ and $\mathcal{F}$.
@@
The inequality in 1 can be interpreted as follows: the "distance" between $\rho$ and $\nu$ is greater when observed from an anchor $\mu$ that differs from $\rho$ and $\nu$. We show that a sufficient condition for Bregman divergences to satisfy this assumption are the following conditions on the Bregman potential $\phi$.
@@
We now discuss the practical implementation of MD on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$ as written in (5). If $\phi_{\mu}$ is pushforward compatible, we have $\nabla\phi_{\mu_{k}}(\mathrm{T}_{k+1})=\nabla_{\mathrm{W}_{2}}\phi((\mathrm{T}_{k+1})_{\#}\mu_{k}\big{)}\circ\mathrm{T}_{k+1}$; but if $\nabla\phi_{\mu_{k}}^{*}$ is unknown, the scheme is implicit in $\mathrm{T}_{k+1}$. A possible solution is to rely on a root finding algorithm such as Newton’s method to find the zero of $\nabla\mathrm{J}$ at each step. However, this procedure may be computationally costly and scale badly w.r.t. the dimension and the number of samples, see Section G.1. Nonetheless, in the special case $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V$ differentiable, strongly convex and $L$-smooth, since $\nabla_{\mathrm{W}_{2}}\mathcal{V}(\mu)=\nabla V$ and $(\nabla V)^{-1}=\nabla V^{*}$, the scheme reads as
$\forall k\geq 0,\ \mathrm{T}_{k+1}=\nabla V^{*}\circ\big{(}\nabla V-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}.$ (9)
This scheme is analogous to MD in $\mathbb{R}^{d}$ [11] and has been introduced as the mirror Wasserstein gradient descent [105]. Moreover, for $V=\frac{1}{2}\|\cdot\|_{2}^{2}$, as observed earlier, we recover the usual Wasserstein gradient descent, i.e. $\mathrm{T}_{k+1}=\mathrm{Id}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})$ [6]. The scheme can also be implemented for Bregman potentials that are not pushforward compatible. For specific $\phi$, it recovers notably (mirrored) SVGD [77, 76, 106] or the Kalman-Wasserstein gradient descent [52]. We refer to Section D.4 for more details.
@@ 4 Preconditioned gradient descent
As seen in Section 2, preconditioned gradient descent on $\mathbb{R}^{d}$ has dual convergence conditions compared to MD. Our goal is to extend these to (1) and $\mathcal{P}_{2}(\mathbb{R}^{d})$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $h:\mathbb{R}^{d}\to\mathbb{R}$ proper and strictly convex on $\mathbb{R}^{d}$. We consider in this section $\phi_{\mu}^{h}(\mathrm{T})=\int h\circ\mathrm{T}\ \mathrm{d}\mu$ and $\mathrm{d}(\mathrm{T},\mathrm{Id})=\phi_{\mu_{k}}^{h}\big{(}(\mathrm{Id}-\mathrm{T})/\tau\big{)}\tau=\int h\big{(}(x-\mathrm{T}(x))/\tau\big{)}\tau\ \mathrm{d}\mu_{k}(x)$. This type of discrepancy is analogous to OT costs with translation-invariant ground cost $c(x,y)=h(x-y)$, which have been popular as they induce an OT map [102, Box 1.12]. Such costs have been introduced e.g. in [36, 65] to promote sparse transport maps. More generally, for $\phi_{\mu}$ strictly convex, proper, differentiable and superlinear, we have $(\nabla\phi_{\mu})^{-1}=\nabla\phi_{\mu}^{*}$ and the following theory is still valid. For simplicity, we leave studying more general $\phi$ for future works. Here, the scheme (1) results in:
@@
Deriving the first order conditions similarly to Section 3, we obtain the following update:
@@
Notice that for $h=\frac{1}{2}\|\cdot\|_{2}^{2}$ the squared Euclidean norm, $\phi_{\mu}^{h}$ and $\phi_{\mu}^{h^{*}}$ recover the squared $L^{2}(\mu)$ norm, and schemes (4) and (10) coincide. The scheme (10) is analogous to preconditioned gradient descent [82, 111, 63], which provides a dual alternative to mirror descent. For the latter, the goal is to find a suitable preconditioner $h^{*}$ allowing to have convergence guarantees, or to speed-up the convergence for ill-conditioned problems. It was recently considered on the Wasserstein space by Cheng et al. [29] and Dong et al. [41] with a focus on the KL divergence as objective $\mathcal{F}$ and for $h=\|\cdot\|_{p}^{p}$ with $p>1$ [29] or $h$ quadratic [41]. Moreover, their theoretical analysis was mostly done using the continuous formulation $\partial_{t}\mu_{t}-\mathrm{div}\big{(}\mu_{t}\nabla h^{*}\circ\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{t})\big{)}=0$ [1], while we focus on deriving conditions for the convergence of the discrete scheme (11) for more general functionals objectives.
@@
Inspired by [82], we now provide a descent lemma on $\big{(}\phi_{\mu_{k}}^{h^{*}}(\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}))\big{)}_{k}$ under a technical inequality between the Bregman divergences of $\phi_{\mu_{k}}^{h^{*}}$ and ~$\tilde{\mathcal{F}}_{\mu_{k}}$ for all $k\geq 0$. Additionally, we also suppose that $\mathcal{F}$ is convex along the curves generated by $\mathrm{T}=\mathrm{Id}$ and $\mathrm{T}_{k+1}$. This last hypothesis ensures that $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})\geq 0$, and thus that $\big{(}\phi_{\mu_{k}}^{h^{*}}(\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}))\big{)}_{k}$ is non-increasing. Analogously to the Euclidean case, $\phi_{\mu}^{h^{*}}$ quantifies the magnitude of the gradient, and provides a second quantifier of convergence leading to possibly different efficient methods compared to mirror descent [63]. The proof relies mainly on the three-point identity (see e.g. [50, Appendix B.7] or Lemma 26) and algebra with the definition of Bregman divergences.
@@
Under an additional assumption of a reverse inequality between the Bregman divergences of $\phi_{\mu_{k}}^{h^{*}}$ and ~$\tilde{\mathcal{F}}_{\mu_{k}}$, and assuming that $\phi_{\mu}^{h^{*}}$ attains its minimum in 0, we can show the convergence of the gradient quantified by $\phi^{h^{*}}$ (see Lemma 19), and the convergence of $\big{(}\mathcal{F}(\mu_{k})\big{)}_{k}$ towards the minimum of $\mathcal{F}$.
@@
Let $\alpha\geq 0$ and $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be the minimizer of $\mathcal{F}$. Assume the conditions of Proposition 4 hold, and for $\bar{\mathrm{T}}=\operatorname*{argmin}_{\mathrm{T},\mathrm{T}_{\#}\mu_{k}=\mu^{*}}\ \mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T}),$ $\alpha\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\bar{\mathrm{T}})\leq\mathrm{d}_{\phi_{\mu_{k}}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\bar{\mathrm{T}}_{\#}\mu_{k})\circ\bar{\mathrm{T}},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}$. Then, for all $k\geq 1$, since $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu^{*})=0$ and $\phi_{\mu_{k}}^{h^{*}}(0)=h^{*}(0)$,
@@
The proofs of Proposition 4 and Proposition 5 can be found in Section H.4 and Section H.5.
We now discuss sufficient conditions to obtain the inequalities between the Bregman divergences required in Proposition 4 and Proposition 5. Maddison et al. [82] showed on $\mathbb{R}^{d}$ for a cost $h$ and an objective function $g$, that these conditions were equivalent to $\beta$-smoothness and $\alpha$-convexity of the preconditioner $h^{*}$ (analogous to $\phi_{\mu}^{*}$) relative to the convex conjugate of the objective $g^{*}$ (analogous to ~$\tilde{\mathcal{F}}_{\mu}^{*}$). To write the inequalities we assumed as a relative smoothness/convexity property of $\phi_{\mu_{k}}^{h^{*}}$ w.r.t. ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$, we would need at least to ensure that ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$ is differentiable, as to define its Bregman divergence according to Definition 1, e.g. by assuming ~$\tilde{\mathcal{F}}_{\mu_{k}}$ strictly convex and superlinear (see Lemma 16). The latter is true for several examples of functionals $\mathcal{F}$ we already mentioned, such as potential or interaction energies with strongly convex potentials. In this case, the inequality between the Bregman divergences in Proposition 4 is equivalent with the smoothness of $\phi^{h^{*}}_{\mu}$ relative to ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\circ\mathrm{T}_{k+1}\big{)}_{\#}\mu_{k}$. In particular, for $\mathcal{F}$ a potential energy, the conditions coincide with those of [82] in $\mathbb{R}^{d}$. We refer to Section E.1 for more details.

### v2

@@ 3 Mirror descent
For every $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let $\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ be strictly convex, proper and differentiable and assume that the (sub)gradient $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\in L^{2}(\mu)$ exists. In this section, we are interested in analyzing the scheme (1) where the cost $\mathrm{d}$ is chosen as a Bregman divergence, i.e. $\mathrm{d}_{\phi_{\mu}}$ as defined in Definition 1. This corresponds to a mirror descent scheme in $\mathcal{P}_{2}(\mathbb{R}^{d})$. For $\tau>0$ and $k\geq 0$, it writes:
@@
Iterates of mirror descent.
In all that follows, we assume that the iterates (4) exist, which is true e.g. for a superlinear $\phi_{\mu_{k}}$, since the objective is a sum of linear functions and of the continuous $\phi_{\mu_{k}}$. In the previous section, we have seen that the second term in the proximal scheme (4) can be interpreted as a linearization of the functional $\mathcal{F}$ at $\mu_{k}$ for Wasserstein (sub)differentiable functionals. Now define for all $\mathrm{T}\in L^{2}(\mu_{k})$, $\mathrm{J}(\mathrm{T})=\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})+\tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}$. Then, deriving the first order conditions of (4) as $\nabla\mathrm{J}(\mathrm{T}_{k+1})=0$, we obtain $\mu_{k}$-a.e.,
@@
Note that for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|^{2}_{L^{2}(\mu)}$, the update (5) translates as $\mathrm{T}_{k+1}=\mathrm{Id}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})$, and our scheme recovers Wasserstein gradient descent [35, 91]. This is analogous to mirror descent recovering GD when the Bregman potential is chosen as the Euclidean squared norm in $\mathbb{R}^{d}$ [12]. We discuss in Section D.2 the continuous formulation of (4), showing it coincides with the gradient flow of the mirror Langevin [3, 130], the limit of the JKO scheme with Bregman groundcosts [104], Information Newton’s flows [126], or Sinkhorn’s flow [41] for specific choices of $\phi$ and $\mathcal{F}$.
@@
The inequality in 1 can be interpreted as follows: the “distance” between $\rho$ and $\nu$ is greater when observed from an anchor $\mu$ that differs from $\rho$ and $\nu$. We demonstrate that Bregman divergences satisfy this assumption under the following conditions on the Bregman potential $\phi$.
@@
We now discuss the practical implementation of MD on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$ as written in (5). If $\phi_{\mu}$ is pushforward compatible, we have $\nabla\phi_{\mu_{k}}(\mathrm{T}_{k+1})=\nabla_{\mathrm{W}_{2}}\phi\big{(}(\mathrm{T}_{k+1})_{\#}\mu_{k}\big{)}\circ\mathrm{T}_{k+1}$; but if $\nabla\phi_{\mu_{k}}^{*}$ is unknown, the scheme is implicit in $\mathrm{T}_{k+1}$. A possible solution is to rely on a root finding algorithm such as Newton’s method to find the zero of $\nabla\mathrm{J}$ at each step, which we use in Section 5 for $\phi_{\mu}^{W}$ as Bregman potential. However, this procedure may be computationally costly and scale badly w.r.t. the dimension and the number of samples, see Section G.1. Nonetheless, in the special case $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V$ differentiable, strongly convex and $L$-smooth, since $\nabla_{\mathrm{W}_{2}}\mathcal{V}(\mu)=\nabla V$ and $(\nabla V)^{-1}=\nabla V^{*}$, the scheme reads as
$\forall k\geq 0,\ \mathrm{T}_{k+1}=\nabla V^{*}\circ\big{(}\nabla V-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)},$ (9)
and can be implemented on particles, i.e. for $\hat{\mu}_{k}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}^{k}}$, $x_{i}^{k+1}=\nabla V^{*}\big{(}\nabla V(x_{i}^{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\hat{\mu}_{k})(x_{i}^{k})\big{)}$ for all $k\geq 0,\ i\in\{1,\dots,n\}$. This scheme is analogous to MD in $\mathbb{R}^{d}$ [12] and has been introduced as the mirror Wasserstein gradient descent [113]. Moreover, for $V=\frac{1}{2}\|\cdot\|_{2}^{2}$, as observed earlier, we recover the usual Wasserstein gradient descent, i.e. $\mathrm{T}_{k+1}=\mathrm{Id}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})$. The scheme can also be implemented for Bregman potentials that are not pushforward compatible. For specific $\phi$, it recovers notably SVGD and its variants [84, 83, 131, 114] or the Kalman-Wasserstein gradient descent [56]. We refer to Section D.4 for more details.
@@ 4 Preconditioned gradient descent
As seen in Section 2, preconditioned gradient descent on $\mathbb{R}^{d}$ has dual convergence conditions compared to mirror descent. Our goal is to extend these to (1) and $\mathcal{P}_{2}(\mathbb{R}^{d})$. Let $\tau>0$, $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $h:\mathbb{R}^{d}\to\mathbb{R}$ proper and strictly convex on $\mathbb{R}^{d}$. We consider in this section $\phi_{\mu}^{h}(\mathrm{T})=\int h\circ\mathrm{T}\ \mathrm{d}\mu$ and $\mathrm{d}(\mathrm{T},\mathrm{Id})=\phi_{\mu_{k}}^{h}\big{(}(\mathrm{Id}-\mathrm{T})/\tau\big{)}\tau=\int h\big{(}(x-\mathrm{T}(x))/\tau\big{)}\tau\ \mathrm{d}\mu_{k}(x)$. This type of discrepancy is analogous to OT costs with translation-invariant ground cost $c(x,y)=h(x-y)$, which have been popular as they induce an OT map [110, Box 1.12]. Such costs have been introduced e.g. in [39, 70] to promote sparse transport maps. More generally, for $\phi_{\mu}$ strictly convex, proper, differentiable and superlinear, we have $(\nabla\phi_{\mu})^{-1}=\nabla\phi_{\mu}^{*}$ and the following theory is still valid. For simplicity, we leave studying more general $\phi$ for future works. Here, the scheme (1) results in:
@@
Deriving the first order conditions similarly to (5) in Section 3, we obtain the following update:
@@
Notice that for $h=\frac{1}{2}\|\cdot\|_{2}^{2}$ the squared Euclidean norm, $\phi_{\mu}^{h}$ and $\phi_{\mu}^{h^{*}}$ recover the squared $L^{2}(\mu)$ norm, and schemes (4) and (10) coincide. The scheme (10) is analogous to preconditioned gradient descent [89, 119, 68, 75, 76], which provides a dual alternative to mirror descent. For the latter, the goal is to find a suitable preconditioner $h^{*}$ allowing to have convergence guarantees, or to speed-up the convergence for ill-conditioned problems. It was recently considered on the Wasserstein space by Cheng et al. [31] and Dong et al. [44] with a focus on the KL divergence as objective $\mathcal{F}$ and for $h=\|\cdot\|_{p}^{p}$ with $p>1$ [31] or $h$ quadratic [44]. Moreover, their theoretical analysis [1] was mostly done using the continuous formulation. Instead we focus on deriving conditions for the convergence of the discrete-time scheme (11) for more general functionals objectives.
@@
Inspired by [89], we now provide a descent lemma on $\big{(}\phi_{\mu_{k}}^{h^{*}}(\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}))\big{)}_{k}$ under a technical inequality between the Bregman divergences of $\phi_{\mu_{k}}^{h^{*}}$ and ~$\tilde{\mathcal{F}}_{\mu_{k}}$ for all $k\geq 0$. Additionally, we also suppose that $\mathcal{F}$ is convex along the curves generated by $\mathrm{S}=\mathrm{T}_{k+1}$ and $\mathrm{T}=\mathrm{Id}$. This last hypothesis ensures that $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})\geq 0$, and thus that $\big{(}\phi_{\mu_{k}}^{h^{*}}(\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}))\big{)}_{k}$ is non-increasing. Analogously to the Euclidean case, $\phi_{\mu}^{h^{*}}$ quantifies the magnitude of the gradient, and provides a second quantifier of convergence leading to possibly different efficient methods compared to mirror descent [68]. The proof relies mainly on the three-point identity (see e.g. [54, Appendix B.7] or Lemma 28) and algebra with the definition of Bregman divergences.
@@
Under an additional assumption of a reverse inequality between the Bregman divergences of $\phi_{\mu_{k}}^{h^{*}}$ and ~$\tilde{\mathcal{F}}_{\mu_{k}}$, and assuming that $\phi_{\mu}^{h^{*}}$ attains its minimum in 0, we can show the convergence of the gradient quantified by $\phi^{h^{*}}_{\mu}$ (see Lemma 21), and the convergence of $\big{(}\mathcal{F}(\mu_{k})\big{)}_{k}$ towards the minimum of $\mathcal{F}$.
@@
Let $\alpha\geq 0$ and $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be the minimizer of $\mathcal{F}$. Assume the conditions of Proposition 5 hold, and that for $\bar{\mathrm{T}}=\operatorname*{argmin}_{\mathrm{T},\mathrm{T}_{\#}\mu_{k}=\mu^{*}}\ \mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T}),\;$ $\alpha\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\bar{\mathrm{T}})\leq\mathrm{d}_{\phi_{\mu_{k}}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\bar{\mathrm{T}}_{\#}\mu_{k})\circ\bar{\mathrm{T}},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}$. Then, for all $k\geq 1$, since $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu^{*})=0$ and $\phi_{\mu_{k}}^{h^{*}}(0)=h^{*}(0)$,
@@
The proofs of Proposition 5 and Proposition 6 can be found respectively in Section H.5 and Section H.6.
We now discuss sufficient conditions to obtain the inequalities between the Bregman divergences required in Proposition 5 and Proposition 6. Maddison et al. [89] showed on $\mathbb{R}^{d}$ for a cost $h$ and an objective function $g$, that these conditions were equivalent to $\beta$-smoothness and $\alpha$-convexity of the preconditioner $h^{*}$ (analogous to $\phi_{\mu}^{*}$) relative to the convex conjugate of the objective $g^{*}$ (analogous to ~$\tilde{\mathcal{F}}_{\mu}^{*}$). To write the inequalities we assumed as a relative smoothness/convexity property of $\phi_{\mu_{k}}^{h^{*}}$ w.r.t. ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$, we would need at least to ensure that ~$\tilde{\mathcal{F}}_{\mu_{k}}^{*}$ is differentiable, in order to define its Bregman divergence according to Definition 1. This can be done e.g. by assuming ~$\tilde{\mathcal{F}}_{\mu_{k}}$ strictly convex and superlinear (see Lemma 18). The latter is true for several examples of functionals $\mathcal{F}$ we already mentioned, such as potential or interaction energies with strongly convex potentials.
Under this assumption, we can show that the inequalities in Proposition 5 and Proposition 6 are implied by relative smoothness and convexity along suitable curves.
Proposition 7.
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{T}\in L^{2}(\mu)$. Assume ~$\tilde{\mathcal{F}}_{\mu}^{*}$ is Gâteaux differentiable and define $\mathcal{F}_{\mu}^{*}$ on $t\mapsto\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ as $\mathcal{F}_{\mu}^{*}(\mu_{t})=\tilde{\mathcal{F}}_{\mu}^{*}(\mathrm{T}_{t})$ for $\mathrm{T}_{t}=(1-t)\mathrm{U}+t\mathrm{S}$ for all $t\in[0,1]$, $\mathrm{S},\mathrm{U}\in L^{2}(\mu)$.
If $\phi^{h^{*}}$ is $\beta$-smooth relative to $\mathcal{F}_{\mu}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T}\big{)}_{\#}\mu$. Then,
$\mathrm{d}_{\phi_{\mu}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\big{)}\leq\beta\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{Id},\mathrm{T}).$ (14)
Likewise, if $\phi^{h^{*}}$ is $\alpha$-convex relative to $\mathcal{F}_{\mu}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T}\big{)}_{\#}\mu$, then
$\alpha\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{Id},\mathrm{T})\leq\mathrm{d}_{\phi_{\mu}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\big{)}.$ (15)
In particular, for $\mathcal{F}$ a potential energy, the conditions coincide with those of [89] in $\mathbb{R}^{d}$. We refer to Section E.1 for more details.